the ellipse used to form a lithotripter has a major axis of 12 ft. and a minor axis of 5 ft. how far from the center are the foci?
Answer by TimothyLamb(4379)

---
f = sqrt( a^2 - b^2 )
a = major radius
b = minor radius
---
f = sqrt( 6^2 - 2.5^2 )
f = sqrt( 29.75 )
f ~= 5.4543561
---
answer:
the distance from the center of the ellipse to either foci is about 5.454 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php